OpenStudy (anonymous):

sin(A+B) + sin(A-B)=2sinA cos B prove identity..

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

Okay i am good at math but this just confuses me

OpenStudy (y1066380):

I found that on the Internet, I hope this can help you. sin(a + b) + sin(a - b) = 2sin(a)cos(b) LHS = sin(a + b) + sin(a - b) using the sum formula for sine LHS = sin(a)cos(b) + cos(a)sin(b) + sin(a)cos(b) - cos(a)sin(b) LHS = sin(a)cos(b) + sin(a)cos(b) + [cos(a)sin(b) - cos(a)sin(b)] LHS = 2sin(a)cos(b) + ( 0 ) LHS = 2sin(a)cos(b) RHS = 2sin(a)cos(b) therefore, LHS = RHS

OpenStudy (anonymous):

@y1066380 , thank you!